Determine deltaE for the process H2O(l) --> H2O(g) at 25 degrees Celsius and 1 atm. H2O(l) deltaHf = -286 H2O(g) deltaHf = -242
Answer #1

Ho = Hof products - Hof reactants

Ho = Hof H2O (g) - Hof H2O (l)

Ho = (-242 kJ/mol) - (286 kJ/mol)

Ho = -242 kJ/mol + 286 kJ/mol

Ho = 44 kJ/mol

Expansion work, wex = (ng) * (R) * (T)

where ng = change in the number of gaseous moles = 1

R = constant = 8.314 J/mol-K

T = temperature = 25 oC = 298 K

wex = (1) * (8.314 J/mol-K) * (298 K)

wex = 2500 J/mol

wex = 2.5 kJ/mol

H = E + wex

E = H - wex

E = (44 kJ/mol) - (2.5 kJ/mol)

E = 41.5 kJ/mol